OpenStudy (anonymous):

The endpoints of segment EF are E(-2, 3) and F(5, -3). Find the coordinates of the midpoint of EF.

hartnn (hartnn):

the formula for co-ordinates of mid-point (m,n) when two points (x1,y1)(x2,y2) given is: m=(x1+x2)/2 n=(y1+y2)/2 does this help ?

OpenStudy (campbell_st):

just average the x values and average the y values to find the mid point

OpenStudy (anonymous):

@hartnn the answers are A.(3, 0),B.(½, 2),C.(1 ½, 0),D.(0, 5)

hartnn (hartnn):

what is the average of 5 and -2 or what is (5-2)/2 ??

OpenStudy (anonymous):

that would be 4

OpenStudy (campbell_st):

nope make sure you include the brackets as Hartnn has shown when you use the calculator

hartnn (hartnn):

nopes. (5-2)/2=3/2=1 1/2 and only one option has x co-ordinate as 1 1/2

OpenStudy (anonymous):

so it would be 1 and one half and 0 1 half

hartnn (hartnn):

i couldn't understand what u said,but only C. has x-coordinate as 1 1/2